# Neuro-Symbolic Embedding for Short and Effective Feature Selection via Autoregressive Generation

### v1

@@ 1. Introduction
Feature selection refers to selecting the optimal feature subset that can greatly enhance downstream models. Effective feature selection offers numerous benefits, such as the elimination of redundant features, conservation of computational resources, acceleration of the model learning and deployment, and enhancement of model performance. This technique has been widely integrated into many research domains, such as bioinformatics, financial analysis, and urban computing, among others (Wang et al., 2024a; Ying et al., 2024).
@@ 2. Problem Statement
We formulate the feature selection task as a sequential generative learning task. Given a dataset $D=\{X,Y\}$, where $X$ is a feature set and $Y$ is the target label. We collect the triplet group $R=\{\mathbf{f}_{i},v_{i},u_{i}\}^{n}_{i=1}$ as training data by applying feature selection method on $D$, where $\mathbf{f}_{i}=[f_{1},f_{2},...,f_{m}]$ is a feature subset token sequence; $v_{i}$ is the performance; $u_{i}$ is the redundancy. In this paper, we preserve the knowledge of feature selection existing in $R$ into a continuous embedding space $\varepsilon$ by jointly optimizing an encoder $\phi$, a decoder $\psi$, a performance evaluator $\vartheta$, and a redundancy evaluator $\kappa$. We will search the best embedding $\mathbf{e}^{*}$ within the learned continuous embedding space and reconstruct the optimal feature subset token sequence $\mathbf{f}^{*}$ by:
@@ 3.2. Data Collection
Control Stage: Each feature agent take an action to select or deselect corresponding feature based on their policy networks. These actions will change the selected feature subset, leading a new environment. Also, an overall reward is generated through this process. To quantify it, we introduce a performance evaluator. Specifically, we adopt random forest as the performance evaluator. Because it is robust, stable, widely used in competitions, and can reduce the prediction accuracy variation caused by downstream models. For each feature subset, we can calculate the downstream task accuracy to most intuitively reflect its performance. In reward assignment, the reward is split equally and then assigned to each participating agent. Non-participating agents receive no reward.
@@ 3.3. Data Augmentation
Since our perspective is to treat feature subset embedding as deep sequential learning steered by encoder, decoder, and evaluators, it raises a challenge that the selected feature subsets are non-ordinal sets, making feature knowledge embedding unclear. In addition, data collection is a time-consuming process. Balancing the diversity of data and time is challenging. To solve this problem, we propose a shuffling-based augmentation method to represent a non-ordinal feature subset as a feature sequence. For instance, a triple data with an performance and redundancy, $\{(f_{1},f_{2},f_{3}),0.83,0.71\}$, can be converted into $\{f_{1}f_{2}f_{3},0.83,0.71\}$. In classic computer vision research, people can use rotation or flipping to augment image data. Similarly, since a feature subset is order-agnostic, we can shuffle the feature sequence $\{f_{1}f_{2}f_{3},0.83,0.71\}$ to generate more semantically equivalent data points: $\{f_{2}f_{1}f_{3},0.83,0.71\}$, $\{f_{3}f_{2}f_{1},0.83,0.71\}$ to augment training data. The shuffling augmentation can increase data sizes to fight data sparsity, improve data diversity, reduce data overfitting, construct an empirical training set that is closer to the true population, and improve embedding accuracy.
@@ 3.4. Feature Subset Embedding Model
(14) $\begin{split}\mathcal{L}_{rec}&=-logP_{\psi}(\mathbf{x}^{\mathbf{D}})\\ &=-\sum^{D}_{d=1}logP_{\psi}(f_{d}|\mathbf{e}^{*},[f_{1},f_{2},...,f_{d-1}])\end{split}$
@@ 3.5. Optimal Embedding Search and Reconstruction
The decoder keeps generating the next feature token in an iterative and autoregressive fashion, until the decoder generates an end token (i.e., $EOS$), for instance, ”$f_{2}f_{6}f_{5}EOS$”, representing the optimal feature subset: $\{f_{2},f_{5},f_{6}\}$.
4. Experiments
We conducted experiments on 16 datasets to answer the following questions: 1) Can our neuro-symbolic feature selection learning method improve downstream tasks? 2) Is variational transformer based configuration better than other encoder-decoder configurations? 3) Is our unsupervised strategy effective and efficient? 4) How does redundancy minimization influence the sizes of generated feature sets? 5) How does our RL data collector and data augmentation method influence the effectiveness of the model? 6) Is our method robust when various machine learning models are used as a downstream task? 7) How does the time and space complexity of our method change over an increasing feature size? 8) Is our model robust to noises?
@@ 4.1. Experimental Setup
Data Description: We collected 16 datasets, each of which represents one application domain, from UCIrvine, and OpenML. We evaluated our method and baseline methods on two widely-used predictive tasks in machine learning: 1) classification (C); 2) regression (R). Table 1 shows the key statistics of the data sets, including predictive task type, number of instances, number of features.
Evaluation Design: We aims to evaluate whether our AI-enabled neuro-symbolic feature selection method can improve a downstream predictive model. We used Random Forest as our downstream model. This is because Random Forest is an ensemble learning model widely used in many competitions. Random Forest is stable and robust. This allows us to control variations and uncertainty caused by a downstream predictive model, and better observe the performances of our feature selection learning model. We used the Relative Absolute Error (RAE) to measure the performance of a selected feature subset. A higher RAE score indicates a higher quality subset of features. For each of the 16 domain data sets, we randomly split the data set into two subsets (A and B). A was seen by our method. We used A to collect training data of selected features, performance and redundancy(e.g. $\{(f_{1},f_{2},f_{3}),0.83,0.71\}$). In each training data, the selected features were regarded as a symbolic token sequence, along with the performance and redundancy score. Such symbolic token representations were regarded as feature selection knowledge base. Our generative model learned the knowledge about feature selection from such a neuro-symbolic format. B was not seen by our method. After the best feature subset was identified (e.g., $f_{2}f_{5}f_{6}$) by learning from A, we applied the identified feature subset to B in order to evaluate our selected features.
@@
Data Collector: We used a reinforcement learning agents to explore 300 epochs to collect all the possibilities of feature ID token symbols as training data, and randomly shuffled each feature sequence 25 times to augment the training data.
Feature Subset Embedding: toke sequences to vectors: We mapped feature tokens to a 64-dimensional embedding, and used a 2-layer network for both encoder and decoder, with a multi-head setting of 8 and a feed-forward layer dimension of 256. The latent dimensionality of the variational autoencoder was set to 64. The feature subset utility estimator consists of a 2-layer feed-forward network, in which each layer was of 200 dimensions. The values of $\alpha$, $\beta$, and $\gamma$ were 0.8, 0.2, and 0.001, respectively. When we calculated redundancy, the values of $\alpha$, $\beta$, $\gamma$ and $\delta$ were 0.5, 0.3, 0.001 and 0.2. We set the batch size to 64 and the learning rate to 0.0001.
@@
Baseline Algorithms for Overall Comparisons: We compared our method with 9 widely used feature selection algorithms: 1) K-BEST (Yang and Pedersen, 1997) selects the top-k features with the highest importance scores; 2) mRMR (Peng et al., 2005) selects a feature subset by maximizing relevance with labels and minimizing feature-feature redundancy; 3) LASSO (Tibshirani, 1996) shrinks the coefficients of useless features to zero by sparsity regularization to select features; 4) RFE (Granitto et al., 2006) recursively deletes the weakest features; 5) GFS (Fan et al., 2021) is a group-based feature selection method via interactive reinforcement learning; 6) MCDM (Hashemi et al., 2022) ensemble feature selection as a Multi-Criteria Decision-Making problem, which uses VIKOR sort algorithm to rank features based on the judgment of multiple feature selection methods; 7) LASSONet (Lemhadri et al., 2021) is a neural network with sparsity to encourage the network to use only a subset of input features; 8) SARLFS (Liu et al., 2021) is a simplified version of MARLFS to leverage a single agent to replace multiple agents to decide selection actions of all features; 9) MARLFS (Liu et al., 2019) uses reinforcement learning to create an agent for each feature to learn a policy to select or deselect the corresponding feature, and treat feature redundancy and downstream task performance as rewards.
Variant Design of Our Framework for Internal Comparisons: To justify the impacts of different technical components of our method, we introduced several variants of the proposed model: 1) FSNS removes the variational inference component and solely uses the Transformer to create the feature subset embedding space; 2) ${}^{{}^{\prime}}$ adopts LSTM (Hochreiter and Schmidhuber, 1997) to construct the feature subset embedding space; 3) FSNS- collect training data in an unsupervised manner, using Laplacian Score to quantify the performance of feature subset; 4)FSNS+ is also an unsupervised variant model, but guided by the mean redundancy; 5) FSNSM,C,P denotes the redundancy minimized model using mutual information, covariance, and pearson correlation coefficient respectively.
@@
Dataset Task #Sample #Features Original K-Best mRMR LASSO RFE GFS MCDM LASSONet SARLFS MARLFS FSNS
SpectF C 267 44 75.96 81.41 79.16 75.96 80.80 75.96 79.16 79.16 78.21 75.01 89.78(+8.37%)
SVMGuide3 C 1243 21 77.81 76.54 76.54 75.29 79.43 76.95 80.28 74.87 82.48 78.32 84.11(+1.63%)
German Credit C 1001 24 64.88 65.28 67.63 65.87 67.88 65.51 69.21 58.09 69.24 68.02 72.83(+3.59%)
UCI Credit C 30000 25 80.19 80.48 80.48 77.94 80.16 80.16 74.39 74.11 80.05 80.10 80.78(+0.30%)
SpamBase C 4601 57 92.68 91.47 92.02 91.73 92.13 91.84 90.93 89.08 91.28 91.72 93.02(+0.89%)
Ap_omentum_ovary C 275 10936 66.19 82.80 84.49 82.03 83.19 83.02 84.49 83.02 82.09 84.49 86.52(+2.03%)
Ionosphere C 351 34 92.85 91.38 95.69 86.98 95.69 94.22 91.41 85.58 94.16 95.68 97.10(+1.41%)
Activity C 10299 561 96.17 95.78 95.87 95.34 96.17 96.12 95.97 94.47 95.78 95.83 96.60(+0.43%)
Mice-Protein C 1080 77 74.99 80.11 80.54 78.27 79.14 76.85 74.04 74.99 74.10 80.93 82.41(+1.48%)
Openml-586 R 1000 25 54.95 57.68 57.29 60.67 57.64 45.34 56.90 58.28 55.48 57.71 63.28(+2.61%)
Openml-589 R 1000 25 50.95 54.09 54.03 59.74 54.56 59.74 52.96 54.30 55.07 53.46 61.13(+1.39%)
Openml-607 R 1000 50 51.73 53.54 53.72 58.10 55.14 41.94 54.93 54.25 57.11 55.41 60.89(+2.79%)
Openml-616 R 500 50 15.63 22.69 23.44 28.98 22.56 22.93 22.92 16.34 24.09 23.66 41.81(+12.83%)
Openml-618 R 1000 50 46.89 51.79 51.79 47.41 51.08 52.79 52.23 50.51 50.39 52.45 54.38(+1.93%)
Openml-620 R 1000 25 51.01 57.48 57.20 57.99 58.01 45.17 54.92 58.23 54.56 57.80 62.00(+3.77%)
Openml-637 R 500 50 14.95 21.86 21.88 26.02 15.10 4.26 23.41 3.90 19.45 17.10 39.49(+13.47%)
@@ 4.2.1. A Study of Overall Comparisons
This experiment aims to answer: Can the features selected by our method effectively improve downstream tasks compared with other baseline methods? We evaluated our method and nine baseline algorithms on 16 application datasets over three evaluation metrics. For each application domain data, the objective is to verify whether our method can identify better feature subsets from the data subset A and boost predictive performances on the data subset B. Table 1 shows our method generates better feature subsets than several widely-used baseline algorithms over all the 16 application domains, particularly outperforming the best baseline by an average of 3%. Moreover, none of the baseline algorithms can consistently achieve the second-best performance across all domains. The performances of the baseline models shows the limitations of search based and thresholding based strategies, as well as specific model structural (sparsity) assumptions, making them incapable of generalizing to other domains. In contrast, our neurosymbolic method exhibits better task-agnostic performance and robust performances over domains. This can be explained by our perspective of seeing feature selection as neurosymbolic generative AI; that is, using a large deep sequential learning model to embed complex and mechanism unknown feature knowledge into a large embedding space and then formulate discrete feature selection into a deep generative AI problem under effective and automated continuous optimization.

### v2

@@ 1. Introduction
Feature selection refers to selecting the optimal feature subset that can greatly enhance downstream models. Effective feature selection offers numerous benefits, such as the elimination of redundant features, conservation of computational resources, acceleration of the model learning and deployment, and enhancement of model performance. This technique has been widely integrated into many research domains, such as bioinformatics, financial analysis, and urban computing, among others.
@@
(a) Iterative selection view
(b) Generative view
Figure 1. We view feature selection from the generative perspective and reformulate it as a sequence generation task (b) rather than as an iterative discrete selection process (a).
@@
Our Contributions: 1) Neuro-symbolic generative feature selection Learning: We have redefined the feature selection process by transitioning from a traditional discrete decision-making model to a continuous optimization paradigm. This reformulation enhances the adaptability and efficiency of feature selection methods. 2) Reinforcement learning-based data collector: We introduce a novel RL-based data collection framework to autonomously gathers training data in both supervised and unsupervised settings, significantly reducing dependence on manual data collection. 3) Multi-objective optimization: We incorporate both feature-feature redundancy and downstream task performance as dual objectives in the optimization process. This dual-objective approach uncovers succinct and effective feature subsets, alleviating the training burden and overfitting risk of downstream models, while yielding more robust and reliable results.
@@ 2. Problem Statement
We formulate the feature selection task as a sequential generative learning task. Given a dataset $D=\{X,Y\}$, where $X$ is a feature set and $Y$ is the target label. We collect the triplet group $R=\{\mathbf{f}_{i},v_{i},u_{i}\}^{n}_{i=1}$ as training data by applying feature selection method on $D$, where $\mathbf{f}_{i}=[f_{1},f_{2},...,f_{m}]$ is a feature subset token sequence; $v_{i}$ is the performance; $u_{i}$ is the redundancy. In this paper, we preserve the knowledge of feature selection existing in $R$ into a continuous embedding space $\varepsilon$ by jointly optimizing an encoder $\phi$, a decoder $\psi$, a performance evaluator $\vartheta$, and a redundancy evaluator $\kappa$. We will search the best embedding $\mathbf{E}$ within the learned continuous embedding space and reconstruct the optimal feature subset token sequence $\mathbf{f}^{*}$ by:
@@ 3.2. Data Collection
Control Stage: Each feature agent takes an action to select or deselect corresponding feature based on their policy networks. These actions will change the selected feature subset, leading a new environment. Also, an overall reward is generated through this process. To quantify it, we introduce a performance evaluator. Specifically, we adopt random forest as the performance evaluator. Because it is robust, stable, widely used in competitions, and can reduce the prediction accuracy variation caused by downstream models. For each feature subset, we can calculate the downstream task accuracy to most intuitively reflect its performance. In reward assignment, the reward is split equally and then assigned to each participating agent. Non-participating agents receive no reward.
@@ 3.3. Data Augmentation
Since our perspective is to treat feature subset embedding as deep sequential learning steered by encoder, decoder, and evaluators, it raises a challenge that the selected feature subsets are non-ordinal sets, making feature knowledge embedding unclear. In addition, data collection is a time-consuming process. Balancing the diversity of data and time is challenging. To solve this problem, we propose a shuffling-based augmentation method to represent a non-ordinal feature subset as a feature sequence. For instance, a triple data with an performance and redundancy, $\{(f_{1},f_{2},f_{3}),0.83,0.71\}$, can be converted into “$\{f_{1}f_{2}f_{3},0.83,0.71\}$”. In classic computer vision research, people can use rotation or flipping to augment image data. Similarly, since a feature subset is order-agnostic, we can shuffle the feature sequence “$\{f_{1}f_{2}f_{3},0.83,0.71\}$” to generate more semantically equivalent data points: “$\{f_{2}f_{1}f_{3},0.83,0.71\}$”, “$\{f_{3}f_{2}f_{1},0.83,0.71\}$” to augment training data. The shuffling augmentation can increase data sizes to fight data sparsity, improve data diversity, reduce data overfitting, construct an empirical training set that is closer to the true population, and improve embedding accuracy.
@@ 3.4. Feature Subset Embedding Model
(14) $\begin{split}\mathcal{L}_{rec}&=-logP_{\psi}(\mathbf{x}^{\mathbf{D}}|\mathbf{e}^{*})\\ &=-\sum^{D}_{d=1}logP_{\psi}(f_{d}|\mathbf{e}^{*},[f_{1},f_{2},...,f_{d-1}])\end{split}$
@@ 3.5. Optimal Embedding Search and Reconstruction
The decoder keeps generating the next feature token in an iterative and autoregressive fashion, until the decoder generates an end token (i.e., $EOS$), for instance, “$f_{2}f_{6}f_{5}EOS$”, representing the optimal feature subset: $\{f_{2},f_{5},f_{6}\}$.
4. Experimental Results
We conducted experiments on 19 datasets to answer the following questions: 1) Can our neuro-symbolic feature selection learning method improve downstream tasks? 2) Is variational transformer-based configuration better than other encoder-decoder configurations? 3) Is our unsupervised strategy effective and efficient? 4) How does redundancy minimization influence the sizes of generated feature sets? 5) How does our RL data collector and data augmentation method influence the effectiveness of the model? 6) Is our method robust when various machine learning models are used as a downstream task? 7) How does the time and space complexity of our method change over an increasing feature size? 8) Is our model robust to noises?
@@ 4.1. Experimental Setup
Data Description: We collected 19 datasets, each of which represents one application domain, from UCIrvine, and OpenML. We evaluated our method and baseline methods on two widely-used predictive tasks in machine learning: 1) classification (C); 2) regression (R). Table 1 shows the key statistics of the data sets, including predictive task type, number of instances, and number of features.
Evaluation Design: We aim to evaluate whether our AI-enabled neuro-symbolic feature selection method can improve a downstream predictive model. We used Random Forest as our downstream model. This is because Random Forest is an ensemble learning model widely used in many competitions. Random Forest is stable and robust. This allows us to control variations and uncertainty caused by a downstream predictive model, and better observe the performances of our feature selection learning model. We used the Relative Absolute Error (RAE) to measure the performance of a selected feature subset. A higher RAE score indicates a higher quality subset of features. For each of the 16 domain data sets, we randomly split the data set into two subsets (A and B). A was seen by our method. We used A to collect training data of selected features, performance, and redundancy(e.g. $\{(f_{1},f_{2},f_{3}),0.83,0.71\}$). In each training data, the selected features were regarded as a symbolic token sequence, along with the performance and redundancy score. Such symbolic token representations were regarded as feature selection knowledge base. Our generative model learned the knowledge about feature selection from such a neuro-symbolic format. B was not seen by our method. After the best feature subset was identified (e.g., $f_{2}f_{5}f_{6}$) by learning from A, we applied the identified feature subset to B in order to evaluate our selected features.
@@
Data Collector: We used reinforcement learning agents to explore 300 epochs to collect all the possibilities of feature ID token symbols as training data and randomly shuffled each feature sequence 25 times to augment the training data.
Feature Subset Embedding: token sequences to vectors: We mapped feature tokens to a 64-dimensional embedding, and used a 2-layer network for both encoder and decoder, with a multi-head setting of 8 and a feed-forward layer dimension of 256. The latent dimensionality of the variational autoencoder was set to 64. The feature subset utility estimator consists of a 2-layer feed-forward network, in which each layer was of 200 dimensions. The values of $\alpha$, $\beta$, and $\gamma$ were 0.8, 0.2, and 0.001, respectively. When we calculated redundancy, the values of $\alpha$, $\beta$, $\gamma$ and $\delta$ were 0.5, 0.3, 0.001 and 0.2. We set the batch size to 64 and the learning rate to 0.0001.
@@
Baseline Algorithms for Overall Comparisons: We compared our method with 11 widely used feature selection algorithms: 1) STG (Yamada et al., 2020) selects features based on probabilistic relaxation of the $\ell_{0}$ norm of features; 2) RRA (Seijo-Pardo et al., 2017) gathers different feature sets and integrates them based on statistical sorting distributions. 3) K-BEST (Yang and Pedersen, 1997) selects the top-k features with the highest importance scores; 4) mRMR (Peng et al., 2005) selects a feature subset by maximizing relevance with labels and minimizing feature-feature redundancy; 5) LASSO (Tibshirani, 1996) shrinks the coefficients of useless features to zero by sparsity regularization to select features; 6) RFE (Granitto et al., 2006) recursively deletes the weakest features; 7) GFS (Fan et al., 2021) is a group-based feature selection method via interactive reinforcement learning; 8) MCDM (Hashemi et al., 2022) ensemble feature selection as a Multi-Criteria Decision-Making problem, which uses VIKOR sort algorithm to rank features based on the judgment of multiple feature selection methods; 9) LASSONet (Lemhadri et al., 2021) (LNet) is a neural network with sparsity to encourage the network to use only a subset of input features; 10) SARLFS (Liu et al., 2021) is a simplified version of MARLFS to leverage a single agent to replace multiple agents to decide selection actions of all features; 11) MARLFS (Liu et al., 2019) uses reinforcement learning to create an agent for each feature to learn a policy to select or deselect the corresponding feature, and treat feature redundancy and downstream task performance as rewards.
Variant Design of Our Framework for Internal Comparisons: To justify the impacts of different technical components of our method, we introduced several variants of the proposed model: 1) FSNS removes the variational inference component and solely uses the Transformer to create the feature subset embedding space; 2) FSNS adopts LSTM (Hochreiter and Schmidhuber, 1997) to construct the feature subset embedding space; 3) FSNS- collect training data in an unsupervised manner, using Laplacian Score to quantify the performance of feature subset; 4)FSNS+ is also an unsupervised variant model, but guided by the mean redundancy; 5) FSNSM,C,P denotes the redundancy minimized model using mutual information, covariance, and Pearson correlation coefficient respectively.
@@
Dataset Task #Sample #Features Original STG RRA K-Best mRMR LASSO RFE GFS MCDM LNet SARLFS MARLFS FSNS
SpectF C 267 44 75.96 75.01 79.16 81.41 79.16 75.96 80.80 75.96 79.16 79.16 78.21 75.01 89.78(+8.37%)
SVMGuide3 C 1243 21 77.81 77.06 80.28 76.54 76.54 75.29 79.43 76.95 80.28 74.87 82.48 78.32 84.11(+1.63%)
German Credit C 1001 24 64.88 63.40 67.07 65.28 67.63 65.87 67.88 65.51 69.21 58.09 69.24 68.02 72.83(+3.59%)
UCI Credit C 30000 25 80.19 79.96 74.88 80.48 80.48 77.94 80.16 80.16 74.39 74.11 80.05 80.10 80.78(+0.30%)
SpamBase C 4601 57 92.68 92.03 91.39 91.47 92.02 91.73 92.13 91.84 90.93 89.08 91.28 91.72 93.23(+1.10%)
Ap_omentum_ovary C 275 10936 66.19 82.80 84.49 82.80 84.49 82.03 83.19 83.02 84.49 83.02 82.09 84.49 86.52(+2.03%)
Ionosphere C 351 34 92.85 88.35 85.93 91.38 95.69 86.98 95.69 94.22 91.41 85.58 94.16 95.68 97.10(+1.41%)
Activity C 10299 561 96.17 95.39 95.39 95.78 95.87 95.34 96.17 96.12 95.97 94.47 95.78 95.83 96.60(+0.43%)
Mice-Protein C 1080 77 74.99 76.42 74.53 80.11 80.54 78.27 79.14 76.85 74.04 74.99 74.10 80.93 82.41(+1.48%)
Amazon Employee C 32769 9 93.02 93.49 93.09 93.50 93.50 93.23 93.01 91.56 92.87 93.46 93.46 93.49 93.55(+0.05%)
Higgs Boson C 50000 28 69.57 66.88 68.95 69.90 69.90 69.84 69.66 69.71 68.95 68.72 68.32 69.46 70.32(+0.42%)
Openml-586 R 1000 25 54.95 13.17 58.67 57.68 57.29 60.67 57.64 45.34 56.90 58.28 55.48 57.71 63.28(+2.61%)
Openml-589 R 1000 25 50.95 49.59 52.02 54.09 54.03 59.74 54.56 59.74 52.96 54.30 55.07 53.46 61.13(+1.39%)
Openml-607 R 1000 50 51.73 38.43 55.93 53.54 53.72 58.10 55.14 41.94 54.93 54.25 57.11 55.41 60.89(+2.79%)
Openml-616 R 500 50 15.63 16.57 23.83 22.69 23.44 28.98 22.56 22.93 22.92 16.34 24.09 23.66 41.81(+12.83%)
Openml-618 R 1000 50 46.89 3.84 45.36 51.79 51.79 47.41 51.08 52.79 52.23 50.51 50.39 52.45 54.38(+1.93%)
Openml-620 R 1000 25 51.01 32.29 54.74 57.48 57.20 57.99 58.01 45.17 54.92 58.23 54.56 57.80 62.00(+3.77%)
Openml-637 R 500 50 14.95 -3.13 20.91 21.86 21.88 26.02 15.10 4.26 23.41 3.90 19.45 17.10 39.49(+13.47%)
Housing Boston R 506 13 42.13 41.87 42.09 44.69 40.20 30.29 42.65 42.15 41.56 43.01 35.16 39.70 46.33(+1.64%)
@@ 4.2.1. A Study of Overall Comparisons
This experiment aims to answer: Can the features selected by our method effectively improve downstream tasks compared with other baseline methods? We evaluated our method and nine baseline algorithms on 19 application datasets over three evaluation metrics. For each application domain data, the objective is to verify whether our method can identify better feature subsets from the data subset A and boost predictive performances on the data subset B. Table 1 shows our method generates better feature subsets than several widely-used baseline algorithms over all the 19 application domains, particularly outperforming the best baseline by an average of 3%. Moreover, none of the baseline algorithms can consistently achieve the second-best performance across all domains. The performances of the baseline models shows the limitations of search based and thresholding based strategies, as well as specific model structural (sparsity) assumptions, making them incapable of generalizing to other domains. In contrast, our neuro-symbolic method exhibits better task-agnostic performance and robust performances over domains. This can be explained by our perspective of seeing feature selection as neuro-symbolic generative AI; that is, using a large deep sequential learning model to embed complex and mechanism unknown feature knowledge into a large embedding space and then formulate discrete feature selection into a deep generative AI problem under effective and automated continuous optimization.
(a) UCI Credit
(b) SpamBase
(c) Activity
(d) German Credit
(e) Openml-586
(f) Openml-589
(g) Openml-607
(h) Openml-620
Figure 3. Results of different encoder-decoder configurations for downstream tasks, where FSNS adopts LSTM as the backbone, FSNS replaces LSTM with transformer, and FSNS combines VAE and transformer.